# Dance Any Beat: Blending Beats with Visuals in Dance Video Generation

### v1

@@ 4.2 Evaluation Metrics
To assess the quality of generated videos, we use the Fréchet Video Distance (FVD) [50]. For quantitative assessment of image-level quality, we employ SSIM [56], PSNR [15], and LPIPS [67]. Additionally, we use the CLIP-Score (CS) [42] to measure the cosine similarity between CLIP embeddings of the dance video and Wav2CLIP [58] embeddings of the music, evaluating the coherence between the dance video and the music. To evaluate synchronization between the dancer’s movements and the music beat, we introduce the 2D motion-music alignment score (2D-MM Align). This metric, inspired by the Beat Alignment Score [27] for 3D scenes, adapts the concept for 2D environments. The alignment score is calculated as the average distance between each kinematic beat and its closest music beat. 2D-MM Align is defined as:
@@
Figure 6: Exemplar videos generated using our DabFusion. Showcasing dance videos with unseen individuals in various poses. Following preprocessing of the input image, DabFusion demonstrates its ability to produce dance videos of high quality.
[…]
@@
[…]
@@
4.3 Result Analysis
@@
A notable characteristic of the AIST++ dance dataset is its use of nine cameras C01-C09 to capture dance performances, offering multiple perspectives. We analyze four specific angles, summarized in Table 2, to assess their impact on video quality. MM-Diffusion, trained on a subset of the AIST++ dataset using only data from the C01 camera, is used to generate an equivalent number of videos for direct comparison. Our models outperform MM-Diffusion in terms of FVD and SSIM scores, especially for camera position C01, demonstrating superior video quality. Analysis of scores among different cameras indicates that direct camera angles, such as C01 and C03, where the cameras point straight at the dancer, result in higher-quality video generation compared to oblique angles like C02. Historically, I2V research has not emphasized the impact of camera positioning. Our findings reveal that camera placement significantly influences the quality of resulting videos, underscoring its importance in future I2V research. Additionally, we explored the impact of camera distance by comparing videos from C01 and C09, which share the same angle but differ in their proximity to the dancer. The findings suggest that cameras positioned further from the dancer tend to yield better results.
Table 3: Alignment results between our models: video-audio synchrony and motion-beat alignment.
[…]
GT 0.153 0.223 0.152
[…]
@@
GT 0.138 0.231 0.226
[…]
GT 0.063 0.198 0.101
[…]
GT 0.121 0.218 0.134
[…]
Alignment Evaluation. In Table 3, we assess the synchronicity between music-video and music-motion pairings. The results indicate that our enhanced model surpasses the baseline in all evaluation metrics, demonstrating the positive impact of beat information on both video-audio and motion-beat alignment. Figure 5 illustrates the differences, showing that frames generated by our enhanced model more closely resemble the ground truth videos. We also evaluate the results according to dance style categories corresponding to the music. These assessments, detailed in Table 4, exhibit significant disparities in alignment scores among various dance genres. Notably, Ballet Jazz yields the lowest alignment scores, attributed to its challenging choreography compared to other dance styles. Consequently, the alignment scores reflect the varying levels of difficulty inherent to each dance genre.
@@ 4.4 Choreograph Anyone
DabFusion can choreograph individuals whether they belong to the AIST++ dataset or not, requiring several preprocessing steps. First, we use YOLO [36] for object detection and the Segment Anything Model [24] for segmentation to identify and segment the person from unseen scenarios. We then substitute the individual in the image from the AIST++ dataset with the newly segmented person, creating a combined image. This fused image is set as the starting frame to generate dance videos with musical conditions. We randomly select images featuring diverse individuals in various poses to create dance videos. As illustrated in Figure 6, preprocessing of the input image results in a new image compatible with DabFusion for dance video generation.
Table 5: Ablation study for result comparison between 40-frame video and 80-frame video. The results are from our enhanced model.
@@
Table 6: Ablation study for result comparison between Music embedding and text embedding. The results are from our baseline model.
@@
text 0.136 0.181 0.126 200.54
music 0.138 0.202 0.141 194.44
Table 7: Ablation study for result comparison about arbitrary length video generation. The results are from our enhanced model. We use the last frame of the first generated videos as the starting frame of next video to generate longer videos.
@@ 4.5 Ablation Study
Influence of Video Length. We observe that video length impacts benchmark performance. Our models are trained to generate videos of 40 and 80 frames. Due to training time and computational resource considerations, we designate the 40-frame videos as the baseline. This comparison is detailed in Table 5. The results indicate comparable video quality between the two lengths; however, the 80-frame videos demonstrate superior performance in alignment evaluations. This improvement is primarily attributed to the extended music duration, providing more information for enhancing alignment accuracy.
Text Embedding vs. Music Embedding. Given that CLAP is trained on text-audio pairs, we aim to uncover the type of information it extracts from music, hypothesizing that CLAP primarily identifies dance style cues. To test this, we conduct an experiment using our baseline model, focusing solely on the embedding from CLAP. We substitute the music embedding with one derived from CLAP using the text description of the dance style corresponding to the music. The findings, presented in Table 6, indicate that the impact of different embeddings on FVD and CS scores is minimal. However, videos generated with text-based embeddings exhibit inferior motion-beat alignment. This suggests that CLAP extracts more than just dance style information; it also obtains details crucial for enhancing motion-beat synchronization.
Arbitrary-Length Video Generation. Given the specialty of our task, our model theoretically can generate videos corresponding to any length of music. By using the final frame of one video sequence as the initial frame of the subsequent sequence, we can create videos of arbitrary length. We conducted experiments with four varying lengths of videos generated sequentially using our enhanced model; the results are presented in Table 7. As video length increases, quality tends to drop rapidly. This decline is primarily because the final frames generated by the model lack the fidelity of actual video frames, becoming more apparent with longer videos. Additionally, alignment scores diminish due to this quality degradation. For instance, employing a pose estimation algorithm to determine keypoints for calculating the 2D-MM Alignment score becomes less accurate with lower video quality, adversely affecting the alignment score.
@@ 5 Conclusion
In this paper, we explore the intersection of two distinct research areas: conditional image-to-video generation and music-to-dance generation. We introduce a novel approach, DabFusion, that directly generates dance videos from a single image and accompanying music. This research investigates the role of music as a conditional input in image-to-video synthesis and sets a foundational benchmark for direct dance video generation in response to musical cues. We further examine the extraction of crucial information for enhancing motion-music beat alignment, leveraging both the large-scale foundational model CLAP and the signal-processing tool Librosa. Through extensive experimentation, we have established a robust baseline for this novel task.

### v2

@@ 4.2 Evaluation Metrics
To assess the quality of generated videos, we use the Fréchet Video Distance (FVD) [49]. For quantitative assessment of image-level quality, we employ SSIM [55], PSNR [14], and LPIPS [66]. Additionally, we use the CLIP-Score (CS) [41] to measure the cosine similarity between CLIP embeddings of the dance video and Wav2CLIP embeddings of the music, evaluating the coherence between the dance video and the music. To evaluate synchronization between the dancer’s movements and the music beat, we introduce the 2D motion-music alignment score (2D-MM Align). This metric, inspired by the Beat Alignment Score [26] for 3D scenes, adapts the concept for 2D environments. The alignment score is calculated as the average distance between each kinematic beat and its closest music beat. 2D-MM Align is defined as:
@@
Figure 6: Exemplar videos generated using our DabFusion, showcasing dance videos with unseen individuals in various poses.
Table 1: Quantitative results between DabFusion and MM-Diffusion.
@@
DabFusion (ours) 193.98 0.033 26.88 0.959
4.3 Result Analysis
Video Quality Evaluation. To evaluate the quality of the videos generated by our DabFusion, we use the music and the first frame from the test set, resulting in 2,106 generated videos. We visually present outputs from our model alongside ground truth in Fig. 5. As shown in Fig. 5, the quality of the videos generated by our model closely approximates that of real videos. Given the novelty of our task and the lack of directly comparable previous works, we benchmark our DabFusion against MM-Diffusion [39], the leading model in unconditional video generation. We select MM-Diffusion because it is trained on a subset of the AIST++ dataset, closely aligning with our training dataset and demonstrating state-of-the-art video quality performance on AIST++. We randomly generate an equivalent number of videos with MM-Diffusion for comparison. A quantitative comparison between our models and MM-Diffusion is provided in Tab. 1. The results reveal that the quality of videos produced by our models closely matches that of MM-Diffusion. These comparisons underscore the competitive quality of videos generated by our approach.
@@ 4.3 Result Analysis
A notable characteristic of the AIST++ dance dataset is its use of nine cameras C01-C09 to capture dance performances, offering multiple perspectives. We analyze four specific angles, summarized in Tab. 2, to assess their impact on video quality. MM-Diffusion, trained on a subset of the AIST++ dataset using only data from the C01 camera, is used to generate an equivalent number of videos for direct comparison. Our models outperform MM-Diffusion in terms of FVD and SSIM scores, especially for camera position C01, demonstrating superior video quality. Analysis of scores among different cameras indicates that direct camera angles, such as C01 and C03, where the cameras point straight at the dancer, result in higher-quality video generation compared to oblique angles like C02. Additionally, we explore the impact of camera distance by comparing videos from C01 and C09, which share the same angle but differ in their proximity to the dancer. The findings suggest that cameras positioned further from the dancer tend to yield better results.
Table 3: Alignment results for DabFusion.
@@
Ground Truth 0.153 0.223 0.152
DabFusion (ours) 0.142 0.215 0.148
@@
Break Ground Truth 0.138 0.231 0.226
DabFusion (ours) 0.131 0.223 0.229
Ballet Jazz Ground Truth 0.063 0.198 0.101
DabFusion (ours) 0.057 0.193 0.098
House Ground Truth 0.121 0.218 0.134
DabFusion (ours) 0.112 0.213 0.128
Alignment Evaluation. In Tab. 3, we assess the synchronicity between music-video and music-motion pairings. The results indicate that videos generated by our DabFusion closely match the ground-truth videos in terms of alignment metrics. We also evaluate the results according to dance style categories corresponding to the music. These assessments, detailed in Tab. 4, exhibit significant disparities in alignment scores among various dance genres. Notably, Ballet Jazz yields the lowest alignment scores, attributed to its challenging choreography compared to other dance styles. Consequently, the alignment scores reflect the varying levels of difficulty inherent to each dance genre.
@@ 4.4 Choreograph Anyone
DabFusion can choreograph individuals whether they belong to the AIST++ dataset or not, requiring several preprocessing steps. First, we use YOLO [35] for object detection and the Segment Anything Model [23] for segmentation to identify and segment the person from unseen scenarios. We then substitute the individual in the image from the AIST++ dataset with the newly segmented person, creating a combined image. This fused image is set as the starting frame to generate dance videos with musical conditions. We randomly select images featuring diverse individuals in various poses to create dance videos. As illustrated in Fig. 6, DabFusion is capable of generating high-quality dance videos featuring previously unseen individuals in a variety of contexts and poses.
Table 5: Ablation study for result comparison between 40-frame video and 80-frame video.
@@
Table 6: Ablation study for result comparison between DabFusion with and without beat information.
@@
DabFusion $\times$ 0.138 0.202 0.141 194.44
$\checkmark$ 0.142 0.215 0.148 193.98
Table 7: Ablation study for result comparison about arbitrary length video generation. We use the last frame of the first generated videos as the starting frame of next video to generate longer videos.
@@ 4.5 Ablation Study
Influence of Video Length. We observe that video length impacts benchmark performance. Our model is trained to generate videos of 40 and 80 frames. Due to training time and computational resource considerations, we designate the 40-frame videos as the baseline. This comparison is detailed in Tab. 5. The results indicate comparable video quality between the two lengths; however, the 80-frame videos demonstrate superior performance in alignment evaluations. This improvement is primarily attributed to the extended music duration, providing more information for enhancing alignment accuracy.
Influence of Beat Information. With the emergence of large foundation models capable of generating music representation, many current approaches rely exclusively on these models to encode music, often overlooking information from traditional audio analysis tools. In our experiment, we explore the impact of incorporating beat information into our dance video generation process. The findings, presented in Tab. 6, show that while the inclusion of beat information does not substantially alter the overall quality of the generated videos, it does enhance the performance in terms of video-audio synchronization and motion-beat alignment. Specifically, the model that integrates beat information outperforms the one without in all assessed metrics, underscoring the beneficial effects of beat information.
Arbitrary-Length Video Generation. Given the specialty of our task, our model theoretically can generate videos corresponding to any length of music. By using the final frame of one video sequence as the initial frame of the subsequent sequence, we can create videos of arbitrary length. We conduct experiments with four varying lengths of videos generated sequentially using our model; the results are presented in Tab. 7. As video length increases, quality tends to drop rapidly. This decline is primarily because the final frames generated by the model lack the fidelity of actual video frames, becoming more apparent with longer videos. Additionally, alignment scores diminish due to this quality degradation. For instance, employing a pose estimation algorithm to determine keypoints for calculating the 2D-MM Alignment score becomes less accurate with lower video quality, adversely affecting the alignment score.
@@ 5 Conclusion
We introduce the challenge of creating dance videos from music with our model, DabFusion. This task eliminates the need for keypoint annotations, simplifying data use and enhancing dataset compatibility. DabFusion can animate individuals from reference images, aligning their movements to various musical styles. Our specially designed music encoder effectively captures essential features including dance style, movement, and rhythm, vital for dynamic choreography. Using the AIST++ dataset, we demonstrate DabFusion’s ability to produce high-quality dance videos with excellent synchronization between audio and video, setting a strong foundation for further research in dance video generation.